OpenStudy (anonymous):

Solve for x: 1/2 x + 2 = 22.

OpenStudy (anonymous):

1/2x=20 x=20*2/1 x=40

OpenStudy (anonymous):

x is definitely 40 your right leena

OpenStudy (anonymous):

if is 1/(2x+2)=22 then, 22(2x+2)=1 44x+44=1 44x=-43 x=-43/44